# A Demonstration of Slowed Electron Drift for PTOLEMY

### v2

@@ 3.2 Waveform Analysis
Waveform information stored by the PicoScope for triggered pulses is used to filter spurious events (including noise chirps or inverted pulses due to unstable electrical ground) from the data. We apply lenient ( 98% efficient for most data runs) selection criteria to multiple regimes of the 200 s time domain saved for each event (100 s before and after the pulse trigger). These criteria applied to the mean () and standard deviation () of the waveform during various selected time domains are outlined in Table 1. Examples of waveforms that pass and do not pass such requirements are shown in Figures 16. For passing waveforms, the amplitude of the pulse height is proportional to the energy deposition of the impinging particle.
@@ 3.3 Modes of Operation
Two methods are presented to probe characteristics of the slow drift demonstrator: symmetric slow drift and voltage scan (both shown in Figure 18). Symmetric slow drift studies the change in event rate as the magnitude of the central drift plate voltages is symmetrically reduced (keeping a potential of 0 V at the midplane). Using CST, the top row of Figure 19 shows that potential lines disperse outward as the center potential difference decreases (displayed is a cross-section of the center plane in the demonstrator, though is representative of the field in general). With fewer potential lines remaining continuous, lower total event rates are expected for smaller central drift settings ( does no work, necessitating the guiding center of an emitted electron to follow potential lines of the orthogonal field).
Alternatively, a voltage scan maintains a constant potential difference in the central plates (for instance a kV ‘window’) but alters the midplane potential (central plate settings of 1 and 1 kV, 2 and 0 kV, 3 and 1 kV, etc.). As seen in the bottom row of Figure 19, only certain potential lines remain continuous depending on the midplane potential. The angled source disk — which more frequently emits particles pointed towards the diode than away — transmits electrons more often with guiding centers in the upper half of the field cage. A maximal event rate is thus expected at a negative midplane voltage. Results are presented for both 1 kV and 2 kV window voltage scans.
@@ 4.1 Symmetric Slow Drift
Event rates, including error bars from Poisson statistics, of successfully transported electrons are shown in Figure 20 for various symmetric central drift electrode voltage settings. Settings are reported according to the voltage magnitude set to both central drift plates (i.e., 6 kV corresponds to lower and upper plates set at kV, etc.). Central electrode settings with a larger potential difference have a higher event rate for nearly every voltage bin and longer high voltage tail. To study the total event rate for each potential setting, we sum values in all bins above 45 mV (a threshold well above the noise level of the system and consistent with the PIN diode calibration using the beta spectrometer) in Figure 21. A fractional uncertainty of 16% (applied also to window voltage scans) was determined via repeated measurements under slightly altered conditions (noise level, magnetic field strength, chamber orientation, etc.). The distribution exhibits regimes of steep and plateauing event rate as a function of the center potential difference (and a maximal event rate of roughly 10 events per second at kV). Notably, potential differences approaching zero maintain a nonzero event rate, suggesting the residual field produced by the outer/bounce plates facilitates drift for small central plate settings.
@@ 4.2 Window Voltage Scans
Similarly, cumulative event rates are tabulated and plotted as a function of the midplane voltage setting (i.e., average of top and bottom center plates) for both kV and kV window scans in Figure 22. Both distributions are roughly centered around 1 kV and event rates of the 1 kV scan are chiefly smaller than their 2 kV scan counterparts. In Section 5.2, we compare these results with CST counterparts to verify the accuracy of the simulated model.
@@ 5.2 Comparison to Data
Cumulative event rates obtained from CST are presented in Figure 26, exhibiting slightly higher values than their Section 4 counterparts. We demonstrate a correspondence between the two trios of distributions in Figure 27, fitting according to least squares for a scaling factor and horizontal offset (for the 1 kV and 2 kV windows). Event rates from CST are interpolated, and a 1 confidence interval is tabulated accounting for Poisson statistical fluctuation and deviations in event rates from adjusting the acceptance window by mm. A more rigorous study of systematic uncertainties would be required to fit according to likelihood or another uncertainty-based metric. The horizontal offset of 0.11 kV is an indication that the source disk is not perfectly centered with respect to the drift electrodes, which also explains the lower event rates in data compared to simulation for small symmetric settings in Figure 27. The scaling factor of 1.34 on the detected event rate serves as a general approximation to demonstrate the validity of the simulated model. The 25% discrepancy in event rate is likely due to a confluence of reasons including system alignment, impedance from Teflon structures, smaller-than-expected diode active region, etc. In the following section, we proceed with time of flight estimation.
@@ 5.3 Time of Flight Estimation
To evaluate the time of flight of particles in the slow drift region of the demonstrator, we export trajectory information from CST under the largest ( 6 kV) and smallest ( V) central drift settings studied. The guiding center at each point is tabulated by averaging the electron location over one cyclotron period before/after the point of interest.333In the relativistic regime, the period in which a particle completes one full orbit due to cyclotron motion is given as , where is mass, is the Lorentz factor, is charge, and is magnetic field strength. Time of flight in the slow drift region is then determined based on how long the guiding center is situated between the central plates. Figure 28 shows that a 70 keV electron can be slowed to less than of the maximal drift velocity, being trapped for over 60 ns in the center plate. The drift speed observed in the trajectory is consistent with as evaluated from the simulated fields at each point. Deviations in time of flight occur due to the amount of time a given trajectory spends near the bouncing plates (those that bounce more and reflect nearer to the bouncing plate experience a weaker central drift field on average than those that bounce fewer times and farther away, therefore having a slightly longer drift time). Simulations with various particle energies confirm the slow drift time of flight largely falls within the range of 54-66 ns, only decreasing slightly in the case of 130 keV electrons (at which point the cyclotron radius is so large that the electron spends a considerable amount of time under the influence of the outer plate potentials even when its guiding center is situated near the middle). Additional slow drift trajectories are shown in Figure 29. As the electric field produced in the center region is heavily impacted by potentials set on the outer plates for this setup, increasing the dimensions of the slow drift region, thereby reducing fringe field effects, would allow for a slow drift time of flight on the order of s and largely reduce variations in a given trajectory’s time of flight.
@@ 6 Outlook
The control of -decay electron drift has been exhibited. Under current operating conditions, we can manipulate drift speed by a factor of 5 (containing electrons for over 60 ns), though would theoretically be able to increase this by orders of magnitude by altering the dimensions of the field cage design. Serving as a demonstration for the RF region of the PTOLEMY detector, future studies will look to incorporate electron injection from a low magnetic field region and instrument an antenna array to perform CRES measurements akin to the studies performed by the Project 8 Collaboration [33] and under development for PTOLEMY [28, 34].

### v6

@@ 3.2 Waveform Analysis
Waveform information stored by the PicoScope for triggered pulses is used to filter spurious events (including noise chirps or inverted pulses due to unstable electrical ground) from the data. We apply lenient ( 98% efficient for most data runs) selection criteria to multiple regimes of the 200 s time domain saved for each event (100 s before and after the pulse trigger). These criteria applied to the mean () and standard deviation () of the waveform during various selected time domains are outlined in Table 1. The general motivation for the requirements is that triggered noise characteristically has either displaced directly preceding or following the trigger at 100 s, or has extremely large/small in the later time steps. These specific ranges and values were determined from fine-tuning the denoising of multiple waveform subsets. Examples of waveforms that pass and do not pass such requirements are shown in Figures 16. For passing waveforms, the amplitude of the pulse height is proportional to the energy deposition of the impinging particle.
@@ 3.3 Modes of Operation
In order to create multiple, coherent datasets to compare with simulation, we record event rates with the slow drift demonstrator according to two operational modes: symmetric slow drift and voltage scan (both shown in Figure 18). With direct application to PTOLEMY, symmetric slow drift studies the change in event rate as the magnitude of the central drift plate voltages is symmetrically reduced (keeping a potential of 0 V at the midplane). This corresponds to a slowed drift speed in the central region. Using CST, the top row of Figure 19 shows that potential lines disperse outward as the center potential difference decreases (displayed is a cross-section of the center plane in the demonstrator, though is representative of the field in general). With fewer potential lines remaining continuous, lower total event rates are expected for smaller central drift settings ( does no work, necessitating the guiding center of an emitted electron to follow potential lines of the orthogonal field). Symmetric slow drift data on its own, however, would not make an airtight case for system function. An observed reduction in event rates could simply be a result of a lower voltage (less noisy) operational environment, but a correspondence with simulation would nonetheless be seen.
Alternatively, a voltage scan maintains a constant potential difference in the central plates (for instance a kV ‘window’) but alters the midplane potential (central plate settings of 1 and 1 kV, 2 and 0 kV, 3 and 1 kV, etc.). As seen in the bottom row of Figure 19, only certain potential lines remain continuous depending on the midplane potential. A scan of this nature with a stronger magnetic field could be used to spatially profile the activity of a radioactive source, with resolution defined according to window size. For our setup, the angled source disk — which more frequently emits particles pointed towards the diode than away — transmits electrons more often with guiding centers in the upper half of the field cage. A maximal event rate is thus expected at a negative midplane voltage. Results are presented for both 1 kV and 2 kV window voltage scans.
@@ 4.1 Symmetric Slow Drift
Event rates, including error bars from Poisson statistics, of successfully transported electrons are shown in Figure 20 for various symmetric central drift electrode voltage settings. Settings are reported according to the voltage magnitude set to both central drift plates (i.e., 6 kV corresponds to lower and upper plates set at kV, etc.). Central electrode settings with a larger potential difference have a higher event rate for nearly every voltage bin and longer high voltage tail. To study the total event rate for each potential setting, we sum values in all bins above 45 mV (a threshold well above the noise level of the system and consistent with the PIN diode calibration using the beta spectrometer) in Figure 21. Intuitively, a smaller center plate voltage magnitude corresponds to a reduction in event rate. The distribution additionally exhibits regimes of steep and plateauing event rate as a function of the center potential difference (and a maximal event rate of roughly 10 events per second at kV). Notably, potential differences approaching zero maintain a nonzero event rate, suggesting the residual field produced by the outer/bounce plates facilitates drift for small central plate settings. These less intuitive features are due to the geometry of the system and are mirrored in simulation as seen in Section 5.2.
@@ 4.2 Window Voltage Scans
Similarly, cumulative event rates are tabulated and plotted as a function of the midplane voltage setting (i.e., average of top and bottom center plates) for both kV and kV window scans in Figure 22. Both distributions are roughly centered around 1 kV and event rates of the 1 kV scan are chiefly smaller than their 2 kV scan counterparts. As with Section 4.1, attributes of these distributions are largely influenced by the geometry of the system (particularly source disk angling). In Section 5.2, we compare these results with CST counterparts to verify the accuracy of the simulated model.
4.3 Uncertainty
We establish a fractional uncertainty of 16% on data points by repeating measurements of nearly all settings. The value corresponds to the minimum fractional width on each point in the original dataset such that at least of the total repeated measurements fall within the respective interval (i.e., assuming the distribution is roughly Gaussian, an interval containing of events corresponds to 1 ). This broadly encompasses systematic uncertainties associated with slight variability in chamber alignment, variations in noise level, electric/magnetic field stability, and deviations in detector response (due to self-heating among other phenomena). It, however, does not account for all systematic uncertainties for the physical system, in particular ignoring source alignment, detector alignment, trajectory impediment from Teflon structures, magnetic field uniformity, changes in detector response for various electromagnetic fields, active region size of the detector, and other parameters which do not change between measurements. We consider these uncertainties small enough to ignore, though emphasize the PTOLEMY detector will be rigorously characterized in this regard. Uncertainty on the relative height between the source and detector is addressed in simulation results.
@@ 5.2 Comparison to Data
Cumulative event rates obtained from CST are presented in Figure 26, exhibiting slightly higher values than their Section 4 counterparts. We demonstrate a correspondence between the two trios of distributions in Figure 27, fitting according to least squares for a scaling factor and horizontal offset (for the 1 kV and 2 kV windows). Event rates from CST are interpolated, and a 1 confidence interval is tabulated accounting for Poisson statistical fluctuation and deviations in event rates from adjusting the acceptance window by mm. A more rigorous study of systematic uncertainties would be required to fit according to likelihood or another uncertainty-based metric. The horizontal offset of 0.11 kV is an indication that the source disk is not perfectly centered with respect to the drift electrodes, which may partially explain the lower event rates in data compared to simulation for small symmetric settings in Figure 27. The scaling factor of 1.34 on the detected event rate serves as a general approximation to demonstrate the validity of the simulated model. The 25% discrepancy in event rate is likely due to a confluence of reasons including system alignment, impedance from Teflon structures, smaller-than-expected diode active region, etc. In the following section, we proceed with time of flight estimation.
@@ 5.3 Time of Flight Estimation
To evaluate the time of flight of particles in the slow drift region of the demonstrator, we export trajectory information from CST under the studied kV and V central drift settings. As the largest and (roughly) smallest attainable central magnitudes, these trajectories correspond to the fastest and slowest possible drift speeds. The guiding center at each point is tabulated by averaging the electron location over one cyclotron period before/after the point of interest.666In the relativistic regime, the period in which a particle completes one full orbit due to cyclotron motion is given as , where is mass, is the Lorentz factor, is charge, and is magnetic field strength. Time of flight in the slow drift region is then determined based on how long the guiding center is situated between the central plates. Figure 28 shows that a 70 keV electron can be maximally slowed to less than of the fastest drift velocity, being trapped for over 60 ns in the center plate. One would naively expect a much slower drift speed for such a proportionally small voltage setting; unfortunately, for small center plate voltage magnitudes, the fringe electric field produced by other plates dominates the central region, yielding a fairly high lower bound on drift speed. The observed drift speed in the trajectory is consistent with as evaluated from the simulated fields at each point. Deviations in time of flight occur due to the amount of time a given trajectory spends near the bouncing plates (those that bounce more and reflect nearer to the bouncing plate experience a weaker central drift field on average than those that bounce fewer times and farther away, therefore having a slightly longer drift time). Simulations with various particle energies confirm the slow drift time of flight largely falls within the range of 54-66 ns, only decreasing slightly in the case of 130 keV electrons (at which point the cyclotron radius is so large that the electron spends a considerable amount of time under the influence of the outer plate potentials even when its guiding center is situated near the middle). Additional slow drift trajectories are shown in Figure 29. As the electric field produced in the center region is heavily impacted by potentials set on the outer plates for this setup, increasing the dimensions of the slow drift region, thereby reducing fringe field effects, would allow for a slow drift time of flight on the order of s and largely reduce variations for a given trajectory.
@@ 6 Outlook
The control of -decay electron drift in a defined volume has been exhibited. Under current operating conditions, we can slow drift speed by a factor of 5 (containing electrons for over 60 ns), though would theoretically be able to increase this by orders of magnitude by altering the dimensions of the field cage design. Manipulation of drift speed to trap electrons as such is an essential feature for PTOLEMY to suppress backgrounds and perform a sensitive differential energy measurement of the tritium endpoint to determine . The method of voltage scanning as presented additionally holds value as a technique to spatially profile the activity of a radioactive source. Serving as a demonstration for the RF region, future extensions of this project will look to incorporate electron injection from a low magnetic field region and instrument an antenna array to perform CRES measurements akin to the studies performed by the Project 8 Collaboration [62] and under development for PTOLEMY [51, 50].